# An Ising Model of Bursts and Fizzles in Intergroup Violence

### v8

@@ 5 Conclusions
Intervals of synchrony in Fig. 11 are indicated by A and B. A lasted 2.4 seconds and involved all ties in the group; B involved 67% of the ties for approximately 2 seconds prior to maximum participation in violence. Finally, the graph shows intervals of deescalatory action.
@@
The 21 groups where only a single member fought
Of the 21 groups where violence was committed by a single member, 13 were dyads. In 11 of these 21 groups, one or few deescalators in the focal group successfully prevented participants from using violence. In 2 groups, members of the opponent group were able to avoid collective violence by hampering focal group members from joining the fight. In 2 dyads, the participants were separated by their opponents and could not adjust their actions to each other any longer, leaving them to either fight on their own or flee. All these outcomes are consistent with the Ising model once interrupted ties and/or unconditional defectors are simulated. In 4 of the 6 remaining groups, opponents carried knives, a machete, a bat, or looked too intimidating even when unarmed, which apparently lowered $S_{0}$ in the focal groups. In the remaining 2 groups, members took turns attacking a fallen victim instead of using violence simultaneously, perhaps confident that they were in control (low uncertainty; high $S_{0}$). Because we took the same $S_{0}$ values for all conditional cooperators, variations of these values are below the resolution of our simple model, though.
Overview of cases and descriptive statistics
Our dataset contains information about the behavior of 406 individuals in 42 violent incidents. The 59 groups and their key characteristics are listed in Table 1. After focal and opponent groups have been analyzed as listed, they flip roles with respect to cooperation and turmoil production in a second batch of the analysis, except when an opponent is solitary instead of a group.
@@
Table 1: Overview of all groups and their sizes; occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of turmoil ($T$) and of synchrony at first violence outbreak; number of unconditional defectors ($N_{D}$). In group 8_queensday, two focal group members disappeared from view already before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence.
Table 2 provides mean values, percentages, standard deviations, and minimum and maximum values for the indicators used in the analysis.
@@ 6 Burst 23 (39) 0 1
Solving the Ising model
The Ising spin glass model can be solved computationally by the Metropolis algorithm (next subsection). First, it is solved analytically by a mean field analysis for $p=0$. The case when $p>0$ is analyzed in the main text.
Mean field analysis
@@
So far, we only dealt with conditional cooperators, but in our empirical study, there are also unconditional defectors in proportion $p>0$. Accordingly, we define $M_{cc}$ as the average spin value of the conditional cooperators and $M_{ud}$ as the average spin value of the unconditional defectors. Note that $M_{ud}=S^{-}$. We assume that the unconditional defectors are homogeneously distributed across the network. Accordingly, the mean field equation becomes
@@
Critical proportion of unconditional defectors
@@
Solutions of this equation become complex if $p>\frac{S_{0}}{\Delta}$, hence $p\leq\frac{S_{0}}{\Delta}$. Graphical illustrations are in Fig. 12. The choice of $C=1$ and $D=1/2$ implies that there is no (burst of) cooperation if $p>1/3$.
@@
Solving the Ising model computationally
Here, the Metropolis algorithm is used [3]. Surprisingly, $p_{c}$ of simulated, fully connected networks is very close to the mean field calculation of $p_{c}$, even in very small networks; see Fig. 13. Given $C=1$ and $D=1/2$, there is no burst if $p_{c}>0.34$. In simulations, $p_{c}$ is proportionally less precise in smaller networks (Table 3), and the prediction for the triad (red dot in Fig. 13) is empirically false, whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which did not happen in the simulations. The critical threshold is nearly density independent: if in (sufficiently large) simulated networks, density is decreased by two orders of magnitude, $p_{c}$ increases from $\approx 0.34$ to $\approx 0.35$. The degree distribution has no effect on $p_{c}$.
Figure 13: For simulated networks with sizes $3\leq n\leq 100$, the minimum number of unconditional defectors who prevent a burst is plotted. The red dot marks the triad (1 defector prevents a burst), which conflicts with the empirical data (1 defector and a burst of 2).

### v9

@@ 10 Analysis of coded and plotted videos
Intervals of synchrony in Fig. S14 are indicated by A and B. A lasted 2.4 seconds and involved all ties in the group; B involved 67% of the ties for approximately 2 seconds prior to maximum participation in violence. Finally, the graph shows intervals of deescalatory action.
@@
11 Overview of cases and descriptive statistics
Our dataset contains information about the behavior of 406 individuals in 42 violent incidents. The 59 groups and their key characteristics are listed in Table SLABEL:table:long. After focal and opponent groups have been analyzed as listed, they flip roles with respect to cooperation and turmoil production in a second batch of the analysis, except when an opponent is solitary instead of a group.
@@ 11 Overview of cases and descriptive statistics
Table SLABEL:table:long: Overview of all groups and their sizes; occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of turmoil ($T$) and of synchrony at first violence outbreak; number of unconditional defectors ($N_{D}$). In group 8_queensday, two focal group members disappeared from view already before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence.
Table S2 provides mean values, percentages, standard deviations, and minimum and maximum values for the indicators used in the analysis.
@@ 6 Burst 23 (39) 0 1
12 Solving the Ising model
The Ising model can be solved computationally by the Metropolis algorithm (main text). Here, it is solved analytically by a mean field analysis, first for $p=0$ and then $p>0$. Subsequently, deviations of the computational results from the mean field result are pointed out.
12.1 Mean field analysis
@@
This result without unconditional defectors was inferred in an earlier study [10]. Here, we also deal with unconditional defectors in proportion $p>0$. Accordingly, we define $M_{cc}$ as the average spin value of the conditional cooperators and $M_{ud}$ as the average spin value of the unconditional defectors. Note that $M_{ud}=S^{-}$. We assume that the unconditional defectors are homogeneously distributed across the network. Accordingly, the mean field equation becomes
@@
12.2 Critical proportion of unconditional defectors
@@
Solutions of this equation become complex if $p>\frac{S_{0}}{\Delta}$, hence $p\leq\frac{S_{0}}{\Delta}$. Graphical illustrations are in Fig. S15. The choice of $C=1$ and $D=1/2$ implies that there is no (burst of) cooperation if $p>1/3$.
@@
12.3 Solving the Ising model computationally
The Ising model was solved computationally, in Fortran for speed and in R for comfort. Surprisingly, $p_{c}$ of simulated, fully connected networks is very close to the mean field calculation of $p_{c}$, even in very small networks. Given $C=1$ and $D=1/2$, there is no burst if $p_{c}>0.34$. The critical threshold is nearly density independent: if in (sufficiently large) simulated networks, density is decreased by two orders of magnitude, $p_{c}$ increases from $\approx 0.34$ to $\approx 0.35$. The degree distribution has no effect on $p_{c}$.
In simulations, $p_{c}$ is less precise in smaller networks (Table S3) and the prediction for the triad is empirically false, whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which did not happen in the simulations.